# Spiking Manifesto

### v1

@@ Abstract
Practically everything computers do is better, faster, and more power-efficient than the brain. For example, a calculator crunches numbers more energy-efficiently than any human. Yet AI models are a thousand times less efficient than the brain. These models use artificial neural networks (ANNs) and require GPUs for the multiplication of huge matrices. In contrast, spiking neural networks (SNNs) of the brain have no matrix multiplication and much smaller energy requirements. This manifesto proposes a framework for thinking about popular AI models in terms of spiking networks and polychronization, and for interpreting spiking activity as nature’s way of implementing look-up tables. This offers a way to convert AI models into a novel type of architecture with the promise of a thousandfold improvement in efficiency. Code is available at https://github.com/izhikevich/SNN
@@ I The manifesto
Llama 3 has artificial neurons, one hundred billion (synaptic) parameters, and consumes tens of watts in inference mode. In contrast, the human brain has spiking neurons, one quadrillion synapses, and consumes only 20 watts. Scaled down 10,000-fold to the size of Llama 3, the brain would require only 2 mW – enough to power it for a month with a single AAA battery! There are two reasons for the power efficiency of the brain:
@@
Efficient encoding: Neuronal activations in ANNs are often interpreted as probabilities of spiking or as instantaneous firing rates. Both interpretations ignore the timing of spikes and miss the opportunity to have a combinatorially explosive representation capacity. In Fig. 1b, we depict 60 neurons, each firing only 1 spike during a certain time period. From an ANN’s point of view, this corresponds to all activations being equal to $1$, that is, to a single feature vector, a single embedding, with activation values $(1,1,\ldots,1)\in\mathbb{R}^{60}$. Changing the order of spikes does not change the activation values. However, from an SNN’s point of view, this is just one pattern out of $60$ factorial possible patterns of different relative spiking orders. 60! is greater than $10^{80}$ – the number of particles in the visible universe. Therefore, a single representation value from an ANN’s point of view has a practically infinite representation capacity from an SNN’s point of view.
@@
High representation capacity requires less memory, and sparse activity requires less processing, putting SNNs in the opposite quadrant from ANNs in the memory vs. processing diagram in Fig. 2. While most SNN research emphasizes sparse activity, this manifesto calls to take advantage of the combinatorially explosive representation capacity necessary to achieve the top-right quadrant.
@@
Figure 3: The distinction between ANNs and the SNN described in this manifesto. The linear capacity $e^{\varepsilon^{2}n}$ to represent $\varepsilon$-orthogonal vectors, given in Exercise 5.1.15 in [3], is dwarfed by the combinatorial capacity $n!$.
@@ II The model
Figure 6: Various spiking patterns in layers 1 and 2 are represented by the vector of latencies (timings) $x,y\in\mathbb{R}^{10}$. Transition from $x$ to $y$ is done through $n_{t}=3$ look-up tables. Each look-up table monitors $n_{c}=3$ pairs of ”anchor” neurons, selected randomly during initialization. The first (red) look-up table monitors $(x_{1},x_{2})$, $(x_{2},x_{3})$, and $(x_{3},x_{5})$. At each step, it evaluates 3 conditions, if $x_{1}>x_{2}$, if $x_{2}>x_{3}$, and if $x_{3}>x_{5}$, each resulting in binary 0 or 11. Concatenating these 3 binary values results in an integer between 0 and $7$, e.g., $011_{\rm binary}=3$ in this figure (indexing starts with $000$). The integer is used as an index into the red table to retrieve the vector of synaptic values. These values are used as inputs to the vector of latencies $y$ in layer 2, resulting in various spiking patterns there.
@@
  • Low processing demand. The conversion from a spiking pattern to a look-up table is just a bit concatenation. A look-up table with $n_{c}$ inputs distinguishes $2^{n_{c}}$ different patterns at no additional processing cost. A LUT with $n_{t}$ look-up tables has the capacity to distinguish $2^{n_{t}n_{c}}$ different spiking patterns.
@@ III-A Surrogate gradient
Differentiating this equation with respect to $x_{m}$ gives zero Jacobian vector $\partial y/\partial x_{m}=0$ unless $m$ equals either $a_{i}$} or $b_{i}$}, in which case
@@ IV-C SNN Transformer
for every Query-Key pair $z_{i}$ and $z_{j}$. Here, the positional encoder $PE_{i-j}\in\mathbb{R}^{p}$ is a learnable vector that encodes the relative distance from $z_{i}$ to $z_{j}$. If $p!>n_{inp}$, then $PE$ provides enough context to encode all relative positions as a pattern of spike latencies.
@@ V-B SNN Transformers
We use the same byte-size token prediction task as above to compare the performance of ANN and SNN transformers in Fig. 11. For the ANN transformer, we use exactly the same parameters as in the original paper [24], i.e., $N=6$ layers, $h=8$ heads per layer, $d_{model}=512$, $d_{ff}=2048$, $d_{k}=64$. The ANN simulation, provided by Taylor Kergan from UCSC, was performed without auxiliary training aids such as dropout, Adam optimization, or RMSNorm. Parameters of SNN transformers are provided in Table III.
@@
Since SNN transformers are built on an architecture fundamentally optimized for efficiency and low memory bandwidth demand, we experimented with reducing the embedding dimension to $n=16$. This configuration still has greater encoding capacity, $16!$, than the linear encoding capacity of $\mathbb{R}^{512}$ – the embedding dimension of ANN transformer, discussed in Sect. VIII-A. To further test the benefits of SNNs, we remove the feedforward layer completely. In this ”attention-only” transformer, plotted as blue curve in the figure, the output from attention module in layer $l$ goes directly to the attention module in layer $l+1$ and the whole architecture is just a stack of $N=6$ blocks in Fig. 12b. Unlike scaled dot-product attention (13) of the ANN transformer, LUTs are inherently nonlinear. Therefore, the SNN attention is all you need to perform both token mixing and channel mixing.

### v2

@@ Abstract
Practically everything computers do is better, faster, and more power-efficient than the brain. For example, a calculator performs numerical computations more energy-efficiently than any human. Yet modern AI models are a thousand times less efficient than the brain. These models rely on larger and larger artificial neural networks (ANNs) to boost their encoding capacity, requiring GPUs to perform large-scale matrix multiplications. In contrast, the brain’s spiking neural networks (SNNs) exhibit factorially explosive encoding capacity and compute through the polychronization of spikes rather than explicit matrix–vector products, resulting in lower energy requirements. This manifesto proposes a paradigm for framing popular AI models in terms of spiking networks and polychronization, and for interpreting spiking activity as nature’s way of implementing look-up tables. This suggests a path toward converting AI models into a novel class of architectures with much smaller size yet combinatorially large encoding capacity, offering the promise of a thousandfold improvement in performance. Code is available at https://github.com/izhikevich/SNN
@@ I The manifesto
Llama 3 has artificial neurons, $10^{11}$ (synaptic) parameters, and consumes tens of watts in inference mode. In contrast, the human brain has spiking neurons, $10^{15}$ synapses, and consumes only 20 watts. Scaled down 10,000-fold to the size of Llama 3, the brain would require only 2 mW – enough to power it for a month with a single AAA battery! There are two reasons for the power efficiency of the brain:
@@
Efficient encoding: Neuronal activations in ANNs are often interpreted as probabilities of spiking or as instantaneous firing rates. Both interpretations ignore the timing of spikes and miss the opportunity to have a factorially explosive representation capacity. In Fig. 1b, we depict 60 neurons, each firing only 1 spike during a certain time period. From an ANN’s point of view, this corresponds to all activations being equal to $1$, that is, to a single feature vector, a single embedding, with activation values $(1,1,\ldots,1)\in\mathbb{R}^{60}$. Changing the order of spikes does not change the activation values. However, from an SNN’s point of view, this is just one pattern out of $60$ factorial possible patterns of different relative spiking orders. 60! is greater than $10^{80}$ – the number of particles in the visible universe. Therefore, a single representation value from an ANN’s point of view has a practically infinite representation capacity from an SNN’s point of view.
@@
High representational capacity reduces memory requirements, and sparse activity reduces computations, placing SNNs in the quadrant opposite to ANNs in the memory–processing diagram in Fig. 2. However, current SNN research rarely goes beyond sparsity. This manifesto calls for fully harnessing the combinatorially explosive representational capacity of SNNs to claim and dominate the top-right quadrant.
@@
Figure 3: The distinction between ANNs and the SNN described in this manifesto. The linear capacity $e^{\varepsilon^{2}n}$ to represent $\varepsilon$-orthogonal vectors [3], is dwarfed by the combinatorial capacity $n!$.
@@ II The model
Figure 6: Various spiking patterns in layers 1 and 2 are represented by the vector of latencies (timings) $x,y\in\mathbb{R}^{10}$. Transition from $x$ to $y$ is done through $n_{t}=3$ look-up tables. Each look-up table monitors $n_{c}=3$ pairs of ”anchor” neurons, selected randomly during initialization. The first (red) look-up table monitors $(x_{1},x_{2})$, $(x_{2},x_{3})$, and $(x_{3},x_{5})$. At each step, it evaluates 3 conditions, if $x_{1}>x_{2}$, if $x_{2}>x_{3}$, and if $x_{3}>x_{5}$, each resulting in binary 0 or 11. Concatenating these 3 binary values results in an integer between 0 and $7$, e.g., $011_{\rm binary}=3$ in this figure (indexing starts with $000$). The integer is used as an index into the red table to retrieve the vector of synaptic values. These synaptic values are used as inputs to the vector of latencies $y$ in layer 2, resulting in various spiking patterns there.
@@
  • Low processing demand. The conversion from a spiking pattern to a look-up table is just a bit concatenation. A look-up table with $n_{c}$ inputs distinguishes $2^{n_{c}}$ different patterns at no additional processing cost. A LUT with $n_{t}$ look-up tables has the capacity to distinguish $2^{n_{t}n_{c}}$ different spiking patterns – a giant number.
@@ III-A Surrogate gradient
Differentiating this equation with respect to $x_{m}$ gives zero Jacobian vector $\partial y/\partial x_{m}=0$ for all $m$, except two values, $m={a_{i}}$ and $m={b_{i}}$. In this case
@@ IV-C SNN Transformer
for every Query-Key pair $z_{i}$ and $z_{j}$. Here, the relative positional encoder $PE_{i-j}\in\mathbb{R}^{p}$ is a learnable vector that encodes the relative distance from $z_{i}$ to $z_{j}$ [25]. If $p!>n_{inp}$, then $PE$ provides enough context to encode all relative positions as a pattern of spike latencies.
@@ V-B SNN Transformers
We use the same byte-size token prediction task as above to compare the performance of ANN and SNN decoder-only (GPT-like) transformers in Fig. 11. For the ANN transformer, we use exactly the same parameters as in the original paper [24], i.e., $N=6$ layers, $h=8$ heads per layer, $d_{model}=512$, $d_{ff}=2048$, $d_{k}=64$. The ANN simulation, provided by Taylor Kergan from UCSC, was performed without auxiliary training, such as dropouts, Adam optimization, or RMSNorm. Parameters of SNN transformers are provided in Table III.
@@
Ablation study. Since SNN transformers are built on an architecture fundamentally optimized for efficiency and low memory bandwidth demand, we experimented with reducing the embedding dimension to $n=16$. This configuration still has greater encoding capacity, $16!$, than the linear encoding capacity of $\mathbb{R}^{512}$ – the embedding dimension of ANN transformer, discussed in Sect. VIII-A. To further test the benefits of SNNs, we remove the feedforward layer completely. In this ”attention-only” transformer, plotted as blue curve in the figure, the output from attention module in layer $l$ goes directly to the attention module in layer $l+1$ and the whole architecture is just a stack of $N=6$ blocks in Fig. 12b. Unlike scaled dot-product attention (13) of the ANN transformer, LUTs are inherently nonlinear. Therefore, the SNN attention is all you need to perform both token mixing and channel mixing.